Find molarity, molality, normality of 12g NaOH per 250ml and density= 1.022g
1
Expert's answer
2018-06-04T06:33:48-0400
n(NaOH) = m/M = 12g/40g/mol = 0.3 mol
molarity = n(NaOH)/V(solution) = 0.3mol/0.25L = 1.2 mol/L
normality = E(NaOH)/V(solution) = 0.3mol/0.25L = 1.2 mol/L
m(solution) = 250mL*1.022g/mL = 255.5 g
m(H2O) = 255.5 - 12 = 243.5 g
molality = 0.3mol/0.2435kg = 1.232 mol/kg